OpenStudy (anonymous):

What is the value of 4 to the power of 6 over 4 to the power of 8?

OpenStudy (anonymous):

what do you think is the answer?

OpenStudy (anonymous):

0.0625

OpenStudy (anonymous):



OpenStudy (anonymous):

i think

OpenStudy (anonymous):

oh i think so

OpenStudy (anonymous):

nope try again

OpenStudy (anonymous):

hmmm ill have to think again

OpenStudy (anonymous):

since the bases are same and you are dividing you have to minus the powers so 4^6 divide by 4^8= 4^(6-8)= 4^-2= 1/(4^2)= 1/16

OpenStudy (anonymous):

oooo

OpenStudy (anonymous):

which is 0.0625 as stated by ulitmatewitherlord

OpenStudy (anonymous):

i thought it was multiplying the powers

OpenStudy (anonymous):

thx again